# Smartify: A Multi-Agent Framework for Automated Vulnerability Detection and Repair in Solidity and Move Smart Contracts

### v1

@@ 4 Smartify System Architecture and Workflow
Smartify operates through a five-agent system designed for automated smart contract vulnerability detection and repair. The system functions as shown in Figure 1.
The Smartify system operates in a five-phase process to automatically repair smart contract code. Firstly, in the Input & Initial Audit phase, the smart contract code, written in either Solidity or Move, is fed into the system. The Auditor, an LLM based on Gemma2 9B, analyzes the code to detect potential vulnerabilities and produces a report detailing its findings. Secondly, during Repair Planning, the Architect receives this vulnerability report and formulates a high-level repair plan that outlines the necessary code modifications to address the identified issues. Thirdly, in Code Generation & Refinement, an LLM called CodeGemma which has been fine-tuned for code generation, and is equipped with Retrieval-Augmented Generation (RAG) capabilities, takes the lead. It utilizes separate Move RAG and Solidity RAG components to provide language-specific context. The Code Generator, part of CodeGemma, uses the repair plan to generate the modified code, selecting the appropriate RAG based on the input language and having the capability to perform Solidity to Move translation when necessary. Subsequently, a Self Refinement process is initiated, and the Refiner component iteratively improves the generated code’s quality, readability, and efficiency. Fourthly, in the Validation phase, the Validator (which is the same agent as the Auditor) performs a final security audit on the refined code to ensure that all identified vulnerabilities have been resolved. Finally, the system outputs the repaired smart contract code.
The process may iterate back to step 3 or 4 if the Validator identifies any issues. Each step plays a vital role in ensuring the accurate and secure repair of smart contract code. The workflow is designed to be efficient and effective, leveraging the strengths of each agent to achieve the desired outcome.
@@ 4.1 Agent Prompting Strategy
The agents within Smartify are driven by carefully crafted prompts that guide their actions and ensure consistent performance. We employ a standardized prompt template, adapted from established practices in LLM-based agent systems. The template is structured as follows:
@@
Each agent in our framework is defined by four key components: the Role, which designates the agent’s specific function (such as Auditor, Architect, or Code Generator); the Task, which outlines the agent’s specific objectives; the Instruction, which provides detailed step-by-step guidance using chain-of-thought reasoning; and the Context, which encompasses all necessary information including input code, audit reports, architectural plans, RAG datastore examples, and inter-agent conversation history.
@@ 4.2 Hardware and Model Fine-tuning
The development and deployment of Smartify leveraged a heterogeneous compute environment, utilizing both high-performance GPUs for computationally intensive tasks and a more resource-efficient setup for inference.
@@ 4.2.1 Fine-tuning Setup
    • Data Preprocessing: Tokenization, normalization, and input-output pair creation ensured data consistency and quality.
@@ 4.2.2 Inference Setup
    • Code Generator and Refiner: These agents utilize a fine-tuned CodeGemma model, initially pre-trained on a limited Move corpus and further instruction-tuned to follow Architect-generated ”recipe” patterns. Fine-tuning on Architect outputs ensured it understood these instructions, and pre-training on a limited Move corpus ensured basic syntax understanding.
@@ 5 Experimental Results and Discussion
We run our experiments as defined in our Section 3.3. We report the results as well as the empirical performance of our models. Through that we will try to answer our Research Questions one by one in this section.
Along with Smartify we have ran the benchmark for the following models.
@@
The models were chosen according to the top 8 models at Hugging Face Big Code Leaderboard [40] at the time of this work, and also adding general-purpose models, which are supposed to be better at reasoning.
@@ 5.1 Solidity
This section presents the evaluation results of various code generation models on the task of repairing vulnerabilities in Solidity smart contracts, specifically focusing on the ”Not So Smart Contracts” dataset from the Trail of Bits GitHub repository. This dataset is a collection of intentionally vulnerable Solidity contracts, designed to test the ability of automated tools to detect and repair common security flaws. It contains a diverse set of vulnerabilities, including reentrancy, integer overflow/underflow, access control issues, and timestamp dependence, among others. The dataset has been publicly available for a significant period, raising the possibility that some or all of its contents might be present in the pre-training data of the evaluated models. We analyze the performance of these models based on two key metrics: the number of vulnerabilities fixed and the average inference time, as summarized in Table 5 and Figure 2. We also introduce our framework, Smartify, and demonstrate its effectiveness in enhancing model performance.
@@
The results reveal significant performance disparities among the evaluated models. Among the pre-trained models for Solidity CodeGemma surprisingly emerges as a top performer, successfully fixing 16 vulnerabilities with a relatively low average inference time of 96.5 seconds. This suggests that CodeGemma possesses a strong ability to understand and rectify code vulnerabilities while maintaining reasonable efficiency. However since most of these Solidty smart contracts were part of open githubs repositories, there can be a strong possibility fo these already being part of the pertaining data. Our proposed framework, Smartify (Gemma2+CodeGemma), achieves comparable performance, also fixing 16 vulnerabilities, albeit with a slightly higher average inference time of 112.3 seconds. This increased time is likely due to its iterative multi-agent process, which enables Smartify to leverage the complementary strengths of Gemma2 and CodeGemma, resulting in robust and reliable fixes.
While Smartify here doesn’t immediately show any benefits over codegemma here, we can notice that the same Smartify framework when applied to llama3.1 without any fine-tuning (unlike the Smartify with codegemma) still gives considerable performance boost over vanilla.
@@
Conversely, models like codellama, codequen, deepseekcoder 33b, and llama3.2 show limited effectiveness, fixing only a small number of vulnerabilities. The poor performance of these models could be attributed to several factors, such as insufficient exposure to Solidity code during pre-training or fine-tuning, or architectures ill-suited for vulnerability repair, which requires a deep understanding of both code syntax and security principles. The exceptionally poor performance of models like starcoder2 (marked with an asterisk *), along with incomplete data for opencoder, suggests potential issues with their training data or a fundamental mismatch between their capabilities and the task’s demands. These models might have been trained on an older version of Solidity or different smart contract security practices than those in the Not-So-Smart-Contracts dataset. Moreover, they might prioritize other aspects of code generation, such as code completion, over security-specific tasks like vulnerability repair.
The public availability of the ”Not So Smart Contracts” dataset raises the question of data contamination. Many evaluated models, especially those trained on large, public code corpora, might have encountered this dataset during pre-training, potentially inflating their performance. However, since CodeGemma and Smartify (Gemma2+codegemma) were specifically fine-tuned for this task, the issue of data contamination is likely less significant.
5.2 Move Code Repair
@@
Move Prover [18] - 2 - - - - 47 15 -
@@
This section analyzes the efficacy of various models in repairing vulnerabilities within Move smart contracts, as detailed in Table 6. The evaluation encompasses eight distinct vulnerability categories: Unchecked Return (UR), Infinite Loop (IL), Unnecessary Boolean (UB), Unused Constant (UC), Unused Private Function (UPF), Unnecessary Type Conversion (UTC), Overflow (Ov), and Precision Loss (PL) following the works of Song et al [14]. The metrics presented in the table represent the number of successfully repaired instances for each vulnerability type, with higher values indicating superior performance. The inference time, measured in seconds, is also provided for each model.
@@
The Smartify models, which leverage a combination of Gemma2 with either codegemma or llama3.1, present an interesting case. Smartify (Gemma2+codegemma) and Smartify (Gemma2+llama3.1) outperform several individual models in multiple categories. This is likely because the specialized models are fine-tuned on the Move-specific dataset. For instance, Smartify (Gemma2+codegemma) achieves the highest number of repairs for the Unchecked Return, Infinite Loop, Unused Boolean, Unused Constant, Unused Private Function, Unnecessary Type Conversion, and Overflow categories, showcasing a substantial improvement over individual models in these areas. However, it is worth mentioning that they also have limitations compared to individual models for certain categories like Precision Loss.
This answers our first two research questions.
RQ1 & RQ2 - Code Understanding and Vuln. Detection Yes. Our empirical analysis with Smartify, especially with using a fine-tuned code-gemma and also using a vanilla pre-trained llama3.1, has shown us the effectiveness of the framework’s ability to understand code. And to capture bad practices leading to vulnerability. Especially for a low-resource code like move. Without significant fine-tuning (in the case of llama3.1).
@@
RQ3 - Code Repair Both for solidity and move, we were able to compare the efficacy of our framework with prior works and can see Smartify outperforms all of the existing code models, even very specialized code models trained on move (opencoder [41]) in generating repair codes for detected vulnerabilities.
Furthermore, Smartify’s efficacy extends even when integrating a non-finetuned model like Llama 3.1. Smartify significantly outperforms Llama 3.2 by fixing 14 vulnerabilities compared to Llama 3.2’s single fix, making its performance comparable with the much larger and computationally intensive Llama 3.3 70b. This demonstrates that Smartify’s architecture can enhance even general-purpose language models for code repair, offering a balance between speed and accuracy. Answering our last query:
RQ4 - Generalization Our implementation of Smartify with both fine-tuned code-gemma and llama3.1 as the second agent gave us the opportunity to run our experiments on both sets of LLMs. And the results show that Smartify is able to significantly boost performance even on non-finetuned models compared to a single model.
Comparative analysis reveals trade-offs between model scale and performance in automated code repair. Larger models, such as deepseekcoder 33b and Llama 3.3 70b, exhibit broader repair capabilities but incur higher computational costs and inference times. Conversely, the Gemma2 27b model demonstrates notable proficiency in addressing Overflow vulnerabilities, albeit with limitations in handling Unnecessary Boolean and Unused Constant compared to Llama 3.3 70b. While Llama 3.3 70b outperforms Smartify in overall repair capability, its significantly slower inference speed poses a challenge for practical deployment. Therefore, for real-world, on-device applications, Smartify (Gemma2+codegemma) presents a compelling solution with its balance of strong accuracy and rapid inference.
Insight: Specialized code models like Starcoder [42], Opencoder [41] and deepseekcoder [43] doesn’t necessarily work well even if it’s a coding specific task. While codemodels like codegemma [44] and codellama [39] are much better at understanding instructions and working on code. This helped Smartify for its understanding and fine-tuning for code repairability.
@@
This research also opens up future research directions of the use of this framework for context-aware test case generation.

### v2

@@ 4 Smartify System Architecture and Workflow
Smartify operates through a five-agent system designed for automated intelligent contract vulnerability detection and repair. The system functions as shown in Figure 1.
The Smartify system operates in a five-phase process to automatically repair smart contract code. Firstly, in the Input & Initial Audit phase, the smart contract code, written in either Solidity or Move, is fed into the system. The Auditor, an LLM based on Gemma2 9B, analyzes the code to detect potential vulnerabilities and produces a report detailing its findings. Secondly, during Repair Planning, the Architect receives this vulnerability report and formulates a high-level repair plan that outlines the necessary code modifications to address the identified issues. Thirdly, in Code Generation & Refinement, an LLM called CodeGemma, which has been fine-tuned for code generation and is equipped with Retrieval-Augmented Generation (RAG) capabilities, takes the lead. It utilizes separate Move RAG and Solidity RAG components to provide language-specific context. The Code Generator, part of CodeGemma, uses the repair plan to generate the modified code, selecting the appropriate RAG based on the input language and able to perform Solidity to Move translation when necessary.
Subsequently, a Self-refinement process is initiated, and the Refiner component iteratively improves the generated code’s quality, readability, and efficiency. Fourthly, in the Validation phase, the Validator (the same agent as the Auditor) performs a final security audit on the refined code to ensure all identified vulnerabilities have been resolved. Finally, the system outputs the repaired smart contract code.
The process may iterate to step 3 or 4 if the Validator identifies any issues. Each step is vital in ensuring the smart contract code’s accurate and secure repair. The workflow is designed to be efficient and effective, leveraging each agent’s strengths to achieve the desired outcome.
@@ 4.1 Agent Prompting Strategy
The agents within Smartify are driven by carefully crafted prompts that guide their actions and ensure consistent performance. We employ a standardized prompt template adapted from established practices in LLM-based agent systems. The template is structured as follows:
@@
Each agent in our framework is defined by four key components: the Role, which designates the agent’s specific function (such as Auditor, Architect, or Code Generator); the Task, which outlines the agent’s particular objectives; the Instruction, which provides detailed step-by-step guidance using chain-of-thought reasoning; and the Context, which encompasses all necessary information including input code, audit reports, architectural plans, RAG datastore examples, and inter-agent conversation history.
@@ 4.2 Hardware and Model Fine-tuning
The development and deployment of Smartify leveraged a heterogeneous computing environment, utilizing high-performance GPUs for computationally intensive tasks and a more resource-efficient setup for inference.
@@ 4.2.1 Fine-tuning Setup
    • Data Pre-processing: Tokenization, normalization, and input-output pair creation ensured data consistency and quality.
@@ 4.2.2 Inference Setup
    • Code Generator and Refiner: These agents utilize a fine-tuned CodeGemma model, initially pre-trained on a limited Move corpus and further instruction-tuned to follow Architect-generated "recipe" patterns. Fine-tuning on Architect outputs ensured it understood these instructions, and pre-training on a limited Move corpus ensured basic syntax understanding.
@@ 5 Experimental Results and Discussion
We run our experiments as defined in Section 3.3. We report the results as well as the empirical performance of our models. Through that, we will try to answer our Research Questions individually in this section.
Along with Smartify, we have run the benchmark for the following models.
@@
The models were chosen according to the top 8 models at Hugging Face Big Code Leaderboard huggingfaceCodeModels at the time of this work, and also adding general-purpose models, which are supposed to be better at reasoning.
@@ 5.1 Solidity
This section presents the evaluation results of various code generation models on repairing vulnerabilities in Solidity smart contracts, specifically focusing on the “Not So Smart Contracts” dataset from the Trail of Bits GitHub repository. This dataset is a collection of intentionally vulnerable Solidity contracts designed to test the ability of automated tools to detect and repair common security flaws. It contains diverse vulnerabilities, including re-entrancy, integer overflow/underflow, access control issues, and timestamp dependence, among others. The dataset has been publicly available for a significant period, raising the possibility that some or all of its contents might be present in the pre-training data of the evaluated models. We analyze the performance of these models based on two key metrics: the number of vulnerabilities fixed and the average inference time, as summarized in Table 5 and Figure 2. We also introduce our framework, Smartify, and demonstrate its effectiveness in enhancing model performance.
@@
The results reveal significant performance disparities among the evaluated models. Among the pre-trained models for Solidity CodeGemma surprisingly emerges as a top performer, successfully fixing 16 vulnerabilities with a relatively low average inference time of 96.5 seconds. This suggests that CodeGemma possesses a strong ability to understand and rectify code vulnerabilities while maintaining reasonable efficiency. However since most of these Solidity smart contracts were part of open Githubs repositories, there can be a strong possibility fo these already being part of the pertaining data. Our proposed framework, Smartify (Gemma2+CodeGemma), achieves comparable performance, also fixing 16 vulnerabilities, albeit with a slightly higher average inference time of 112.3 seconds. This increased time is likely due to its iterative multi-agent process, which enables Smartify to leverage the complementary strengths of Gemma2 and CodeGemma, resulting in robust and reliable fixes.
While Smartify does not immediately show any benefits over CodeGemma here, we can notice that the same Smartify framework when applied to Llama3.1 without any fine-tuning (unlike the Smartify with CodeGemma) still gives considerable performance boost over vanilla.
@@
Conversely, models like CodeLlama, CodeQWen, DeepSeekCoder 33b, and Llama3.2 show limited effectiveness, fixing only a small number of vulnerabilities. The poor performance of these models could be attributed to several factors, such as insufficient exposure to Solidity code during pre-training or fine-tuning or architectures ill-suited for vulnerability repair, which requires a deep understanding of code syntax and security principles. The exceptionally poor performance of models like starcoder2 (marked with an asterisk *), along with incomplete data for opencoder, suggests potential issues with their training data or a fundamental mismatch between their capabilities and the task’s demands. These models might have been trained on an older version of Solidity or different smart contract security practices than those in the Not-So-Smart-Contracts dataset. Moreover, they might prioritize other aspects of code generation, such as code completion, over security-specific tasks like vulnerability repair.
@@
Move Prover dill2022fast - 2 - - - - 47 15 -
@@
The public availability of the "Not So Smart Contracts" dataset raises the question of data contamination. Many evaluated models, especially those trained on large, public code corpora, might have encountered this dataset during pre-training, potentially inflating their performance. However, since CodeGemma and Smartify (Gemma2+CodeGemma) were specifically fine-tuned for this task, the issue of data contamination is likely less significant.
5.2 Move Code Repair
This section analyzes the efficacy of various models in repairing vulnerabilities within Move smart contracts, as detailed in Table 6. The evaluation encompasses eight distinct vulnerability categories: Unchecked Return (UR), Infinite Loop (IL), Unnecessary Boolean (UB), Unused Constant (UC), Unused Private Function (UPF), Unnecessary Type Conversion (UTC), Overflow (Ov), and Precision Loss (PL) following the works of Song et al song2024empirical. The metrics presented in the table represent the number of successfully repaired instances for each vulnerability type, with higher values indicating superior performance. The inference time, measured in seconds, is also provided for each model.
@@ 5.2 Move Code Repair
The Smartify models, which take advantage of a combination of Gemma2 with either CodeGemma or Llama3.1, present an interesting case. Smartify(Gemma2+CodeGemma) and Smartify(Gemma2+Llama3.1) outperform several individual models in multiple categories. This is likely because the specialized models are fine-tuned on the Move-specific dataset. For instance, Smartify (Gemma2+CodeGemma) achieves the highest number of repairs for the Unchecked Return, Infinite Loop, Unused Boolean, Unused Constant, Unused Private Function, Unnecessary Type Conversion, and Overflow categories, showcasing a substantial improvement over individual models in these areas. However, it is worth mentioning that they also have limitations compared to individual models for certain categories like Precision Loss.
RQ1 & RQ2 - Code Understanding and Vuln. Detection Yes. Our empirical analysis with Smartify, especially with using a fine-tuned Code-Gemma and also using vanilla pre-trained Llama3.1, has shown us the effectiveness of the framework’s ability to understand code and capture bad practices leading to vulnerability. Especially for a low-resource code like Move, without significant fine-tuning (in the case of Llama3.1), Smartify outperform the larger and computationally intensive models such as Llama 3.3 70b.
@@
RQ3 & RQ4 - Code Repair Both for Solidity and Move, we were able to compare the efficacy of our framework with prior works. We can see Smartify outperforms all of the existing code models, even very specialized code models trained on Move (OpenCoder huang2024opencoder) in generating repair codes for detected vulnerabilities.
Furthermore, Smartify’s efficacy extends even when integrating a non-finetuned model like Llama 3.1. Smartify significantly outperforms Llama 3.2 by fixing 14 vulnerabilities compared to Llama 3.2’s single fix, making its performance comparable with the much more extensive and computationally intensive Llama 3.3 70b. This demonstrates that Smartify’s architecture can enhance even general-purpose language models for code repair, balancing speed, and accuracy.
RQ5 - Generalization Our implementation of Smartify with both fine-tuned Code-Gemma and Llama3.1 as the second agent allowed us to run our experiments on both sets of LLMs. The results show that Smartify can significantly boost performance even on non-finetuned models compared to a single model.
Comparative analysis reveals trade-offs between model scale and performance in automated code repair. Larger models, such as DeepSeekCoder 33b and Llama 3.3 70b, exhibit broader repair capabilities but incur higher computational costs and inference times. Conversely, the Gemma2 27b model demonstrates notable proficiency in addressing Overflow vulnerabilities, albeit with limitations in handling Unnecessary Boolean and Unused Constant compared to Llama 3.3 70b. While Llama 3.3 70b outperforms Smartify in overall repair capability, its significantly slower inference speed poses a challenge for practical deployment. Therefore, for real-world, on-device applications, Smartify (Gemma2+CodeGemma) presents a compelling solution with its balance of substantial accuracy and rapid inference.
Insight: Specialized code models like StarCoder lozhkov2024starcoder, OpenCoder huang2024opencoder and DeepSeekCoder guo2024deepseek doesn’t necessarily work well even if it’s a coding specific task. While code models like CodeGemma team2024codegemma and CodeLlama roziere2023code are much better at understanding instructions and working on code. This helped Smartify for its understanding and fine-tuning for code repairability.
@@
This research also opens up future research directions on using this framework for context-aware test case generation.
6 Ablation Study
To rigorously validate our architectural design and isolate the contribution of each key component, a comprehensive ablation study was conducted. This study systematically deconstructs the Smartify framework to quantify the impact of its core mechanisms—agent specialization, iterative refinement, and retrieval-augmented generation—on overall repair performance huang2025opencoderopencookbooktoptier; jiang2018contractfuzzer.
6.1 Ablation Configurations
Four distinct configurations of the system were evaluated against the Solidity and Move datasets:
  • Full Smartify: The complete five-agent system, including the RAG module for the Code Generator and the Refiner-Validator iterative feedback loop. This represents our proposed approach.
  • Smartify (No Refinement): A version of the framework where the Refiner and Validator agents are disabled. The output is taken directly from the first pass of the Code Generator. This configuration measures the impact of the iterative self-improvement loop on patch quality.
  • Smartify (No RAG): In this setup, the Code Generator operates without the contextual, few-shot examples provided by the RAG system. It relies solely on its fine-tuned knowledge and the Architect’s plan. This configuration is designed to measure the importance of providing language-specific, in-context examples, especially for the low-resource Move language.
  • Single-Agent Baseline (CoT-style): This configuration replaces the multi-agent system with a single, powerful agent (Gemma2 9B). The agent is given a complex, chained prompt that instructs it to sequentially perform the tasks of auditing, planning, and generating the repair. This mimics the linear reasoning workflow of CoT-based approaches like ContractTinker and directly tests the value of our role-based specialization and delegation architecture.
6.2 Analysis
The performance of each configuration was measured using the Pass@1 and Exploit Mitigation Rate metrics. The results, presented in Table 7, provide clear empirical evidence supporting our architectural choices.
Table 7: Ablation Study of Smartify Components on Solidity and Move Datasets
Solidity Move
Configuration Pass@1 (%) Exploit Mit. (%) Pass@1 (%) Exploit Mit. (%)
Full Smartify 26.7 25.0 48.9 45.7
Smartify (No Refinement) 25.0 20.0 44.6 39.1
Smartify (No RAG) 21.7 18.3 28.3 23.9
Single-Agent Baseline 18.3 15.0 21.7 17.4
The results demonstrate a clear and consistent degradation in performance as key components are removed, confirming the positive contribution of each element of the Smartify architecture.
  • Impact of Iterative Refinement: Removing the refinement loop (No Refinement) causes a noticeable drop in both metrics, particularly the Exploit Mitigation Rate (from 25.0% to 20.0% for Solidity). This indicates that while the initial code generation is often correct, the refinement process is crucial for hardening the patch against exploits and catching subtle regressions that the first pass might miss.
  • Impact of RAG: The removal of the RAG module (No RAG) has a significant negative impact across the board, but its effect is most pronounced for the Move language. The Pass@1 score for Move plummets from 48.9% to 28.3%, a relative decrease of over 42%. This strongly supports our hypothesis that providing in-context, language-specific examples is critical for achieving high performance in low-resource languages where the model’s pre-trained knowledge is limited.
  • Impact of Multi-Agent Architecture: The Single-Agent Baseline, designed to mimic a CoT-style approach, performs the worst of all configurations. Its performance is substantially lower than the Full Smartify framework, particularly on the more complex task of exploit mitigation. This finding provides strong evidence that our role-based, delegative architecture is superior to a monolithic, linear reasoning process. By specializing agents for distinct tasks (analysis vs. planning vs. generation), the system achieves a more robust and effective decomposition of the complex program repair problem.
The ablation study empirically validates the design of Smartify. The multi-agent architecture provides a superior structure for complex problem-solving, the refinement loop is essential for ensuring patch quality and security, and the RAG mechanism is a critical component for adapting the framework to new or low-resource programming languages.
@@ 7 Conclusion
Acknowledgments
This work was partially supported by the Sui Academic Research Award. We gratefully acknowledge the Google Developer Expert AI/ML team for providing cloud compute resources that facilitated this research.
We thank Sam Blackshear of Mysten Labs for his valuable feedback on the idea and early drafts of this paper.